# S-duality in higher-derivative corrections of heterotic supergravity

### v3

@@
This study investigates the consistency of heterotic supergravity, where the $B$-field gauge transformation becomes anomalous due to the Green-Schwarz mechanism, under T-duality transformations. We establish that preserving T-duality necessitates an infinite tower of higher-derivative couplings proportional to $e^{-2\Phi}$. For spacetimes featuring a Killing self-dual circle, these couplings remain immune to quantum corrections. Leveraging S-duality in heterotic/type I string theory, we then identify the corresponding couplings in type I theory. Notably, our analysis reveals that standard S-duality transformations do not involve higher-derivative corrections. Building on this insight, we derive the explicit form of the effective action for type I theory at the $\alpha^{\prime}$ order, adhering to a scheme that omits derivatives of the dilaton.
@@ 1 Introduction
On the other hand, while classical couplings are background independent, quantum corrections to the effective action depend on the background [30]. For backgrounds with one Killing self-dual circle, the aforementioned T-dual invariant couplings do not receive quantum corrections, as the two-derivative couplings are free from such corrections [8, 35]. Since these couplings are immune to quantum corrections in the presence of a Killing self-dual circle, their genus dependence, $e^{-2\Phi}$, is exact, enabling their analysis within the framework of S-duality.
@@
[…]
@@
and no scalar transformation applied to the frame $e_{\mu}{}^{\mu_{1}}$. Consequently, the spin connection $\omega_{\mu\mu_{1}}{}^{\nu_{1}}$, like the YM connection $A_{\mu i}{}^{j}$, remains invariant. As a result, the Lorentz Chern-Simons three-form $\Omega_{\mu\nu\rho}$ also remains invariant under S-duality, i.e.,
@@
This replacement is mandated by the Green-Schwarz mechanism [15], ensuring that T-duality consistently determines the coefficients of an infinite set of higher-derivative couplings at the classical level. These couplings extend the two-derivative terms (4) to T-dual couplings at orders $\alpha^{\prime 0}$, $\alpha^{\prime}$, $\alpha^{\prime 2}$, and beyond, all governed by the same overall factor as the leading-order action (4). Since the two-derivative terms do not receive higher-genus corrections due to kinematic reasons [8, 35], T-duality ensures that the higher-derivative components of the T-dual couplings for the Killing self-dual circle also remain free from genus corrections [^3] A similar set of couplings arises in bosonic string theory. Specifically, bosonic theory includes an infinite number of higher-derivative couplings at orders $\alpha^{\prime}$, $\alpha^{\prime 2}$, and beyond, governed by a single overall factor [37]. However, since this factor is not linked by T-duality to the two-derivative leading-order action, these couplings may not be exact and may receive higher-genus corrections.. It is an exact T-dual multiplet that can be studied under S-duality.
@@
Furthermore, the classical theory includes another distinct set of T-dual couplings at orders $\alpha^{\prime 3}$, $\alpha^{\prime 4}$, and higher, characterized by an overall coefficient of $\zeta(3)$. Unlike the first set, this group is not directly associated with the leading-order action (4). As a result, T-duality does not prevent this collection of T-dual couplings from receiving higher-genus corrections. Similarly, there exist additional sets with coefficients $\zeta(5)$ and higher [38], all of which may also receive higher-genus corrections.
To determine the couplings via T-duality, it is essential to have a spacetime dimension compactified into a circle, where all fields remain independent of the Killing coordinate $y$. Consequently, the results are applicable to spacetimes with one Killing circle. While the classical effective action is background-independent—ensuring that the resulting T-duality invariant couplings are valid across all spacetimes—the quantum corrections to this action are influenced by the background. Thus, the quantum corrections governed by T-duality are applicable exclusively to spacetimes with a single Killing self-dual circle [30].
@@
A similar replacement to (10) exists in Type I theory. However, the corresponding $\alpha^{\prime}H\Omega$ and $\alpha^{\prime 2}\Omega\Omega$ terms do not appear in the usual untwisted sector. Instead, they emerge in the twisted sector of the type I effective action [39], which transforms under T-duality to the twisted sector of the type I’ theory [40]. Unlike in heterotic theory, T-duality alone is insufficient to fully determine all couplings at orders $\alpha^{\prime}$, $\alpha^{\prime 2}$, and higher in both type I and type I’ theories due to inherent dimensional differences: the twisted sector in type I’ theory is inherently 9-dimensional, whereas that in type I theory is 10-dimensional [39]. Nevertheless, S-duality applied to the exact T-dual multiplet of heterotic theory offers a framework to identify such couplings in type I theory.
@@
The corresponding couplings in type I theory manifest at different genera. Using the S-duality transformation (3), it can be shown that the couplings at order $\alpha^{\prime n}$ within the exact T-dual multiplet in heterotic theory transform into couplings at genus $e^{(n+k-2)\Phi}$ in type I theory, where $k$ represents half the number of YM field strengths in the couplings. In this paper, we demonstrate how these couplings can be systematically determined by imposing S-duality on the exact T-dual multiplet in heterotic theory. Specifically, we explicitly derive these couplings in type I theory at the $\alpha^{\prime}$ order, adhering to a scheme that excludes dilaton derivatives. It is important to note that the exact T-dual multiplet in heterotic theory remains free from genus corrections only in spacetimes with a single Killing self-dual circle [30]. Consequently, the couplings derived through S-duality are expected to hold valid exclusively for such spacetimes.
@@
The observation that the exact T-dual multiplet in heterotic theory remains unaffected by quantum corrections in spacetimes with a single Killing self-dual circle has significant implications for the type IIA effective action. These implications arise from the application of S-duality between heterotic theory on the manifold $T^{4}$, which includes the circle, and type IIA theory on the manifold $K3$ [42, 16]. The S-duality correspondence between the 6-dimensional massless NS-NS fields in type IIA and their equivalents in heterotic theory is outlined in [43]:
@@
where $\Phi,G_{\mu\nu},H_{\mu\nu\rho}$ are the six-dimensional NS-NS fields of type IIA theory. In this context, one can use the dimensional reduction of the exact T-dual multiplet on the torus $T^{4}$ to determine the corresponding six-dimensional couplings at all orders of $\alpha^{\prime}$ with the exact genus dependence $e^{-2\Phi^{\prime}}$. Using the aforementioned S-duality transformations, one can identify the corresponding higher-derivative six-dimensional couplings in type IIA theory, which appear at different genera. Applying the S-duality transformation to the exact T-dual multiplet, one finds that the couplings at order $\alpha^{\prime n}$ on $T^{4}$ appear at genus $e^{2(n-1)\Phi}$ in type IIA theory on $K3$. For example, the one-loop couplings in type IIA string theory on $K3$ arise at order $\alpha^{\prime}$. These couplings were first identified in [43] through S-duality transformations and subsequently confirmed by S-matrix calculations. Furthermore, they have been utilized in [44, 45] to investigate $\alpha^{\prime}$ corrections to the dyonic string solution in 6-dimensional supergravity. The two-loop couplings in type IIA on $K3$ are at order $\alpha^{\prime 2}$, which may be found by imposing S-duality on the $\alpha^{\prime 2}$ couplings identified in [25] through T-duality.
@@
On the other hand, the aforementioned S-duality can be utilized to establish a no-go theorem. As previously discussed, aside from the exact T-dual multiplet at orders $\alpha^{\prime 0}$, $\alpha^{\prime}$, $\alpha^{\prime 2}$, and so forth, with an overall coefficient of 1, there exist T-dual multiplets at orders $\alpha^{\prime 3}$, $\alpha^{\prime 4}$, and beyond, characterized by an overall coefficient of $\zeta(3)$, along with an infinite number of T-dual multiplets at higher orders [38]. While the exact T-dual multiplet remains quantum-exact, all other T-dual multiplets are classical and subject to quantum corrections. Therefore, in heterotic theory, only the classical couplings at orders $\alpha^{\prime 0}$, $\alpha^{\prime}$, and $\alpha^{\prime 2}$ remain free from quantum corrections. Since the $T^{4}$ reduction of these classical couplings preserves the number of derivatives in the couplings, it follows that no couplings arise at orders $\alpha^{\prime}$ and $\alpha^{\prime 2}$ at the quantum level, i.e., at genus $e^{2(g-1)\Phi^{\prime}}$ for $g>0$, in the six-dimensional heterotic theory. Using the S-duality transformation, one finds no couplings at order $\alpha^{\prime}$ at genus $e^{-2g\Phi}$ and no couplings at order $\alpha^{\prime 2}$ at genus $e^{-2(g-1)\Phi}$ in the six-dimensional type IIA theory. Based on the permitted genus expansion for the effective action in type IIA theory, it is determined that at order $\alpha^{\prime}$, $g=1$, and at order $\alpha^{\prime 2}$, $g=1,2$. Consequently, there must be no classical couplings at order $\alpha^{\prime}$, and no classical or torus-level couplings at order $\alpha^{\prime 2}$ in the six-dimensional type IIA theory. This conclusion is reinforced by arguments derived from supersymmetry (see, for example, [46]). We will demonstrate that the $K3$ reduction of the classical couplings in type IIA theory at order $\alpha^{\prime 3}$, identified via T-duality [26], produces certain $\alpha^{\prime}$ couplings that vanish upon applying field redefinition.
@@ 2 Heterotic/Type I duality
$\Delta B_{\alpha\beta}(\Phi,G,B)+\Delta B_{\alpha\beta}(-\Phi,e^{-\Phi}G,B)$ $=$ $0$ (13)
@@
$+\frac{1}{48}F_{\alpha\beta ij}F^{\alpha\beta ij}{\tilde{H}}_{\gamma\delta\epsilon}{\tilde{H}}^{\gamma\delta\epsilon}-\frac{1}{4}F^{\alpha\beta ij}{\tilde{H}}_{\beta\gamma\delta}\nabla_{\alpha}F^{\gamma\delta}{}_{ij}\Big{]}+e^{-\Phi}\Big{[}\frac{3}{64}F_{\alpha\beta ij}F^{\alpha\beta ij}R+\frac{1}{16}R^{2}$
$-\frac{1}{16}{\tilde{H}}_{\alpha\beta}{}^{\delta}{\tilde{H}}^{\alpha\beta\gamma}{\tilde{H}}_{\gamma}{}^{\epsilon\varepsilon}{\tilde{H}}_{\delta\epsilon\varepsilon}+\frac{1}{96}{\tilde{H}}_{\alpha\beta\gamma}{\tilde{H}}^{\alpha\beta\gamma}{\tilde{H}}_{\delta\epsilon\varepsilon}{\tilde{H}}^{\delta\epsilon\varepsilon}-\frac{1}{2}F_{\alpha}{}^{\gamma ij}F_{\beta\gamma ij}R^{\alpha\beta}-R_{\alpha\beta}R^{\alpha\beta}$
@@
where ${\tilde{H}}=e^{\Phi}H$. Note that the YM couplings in the first bracket are double traces. The couplings in the last bracket occur at the disk level, while the others appear at higher genus levels. As mentioned in the Introduction section, these couplings are valid for spacetimes with a single Killing self-dual circle because the original couplings in the heterotic theory (15) are exact only in such spacetimes.
At the four-derivative order, the disk-level effective action of type I theory incorporates the single-trace coupling ${\rm Tr}(F^{4})$. In heterotic theory, this coupling emerges as a one-loop contribution and is related, via a supersymmetry transformation, to the anomaly cancellation mechanism in 10-dimensional spacetime. Consequently, it represents an exact coupling in 10 dimensions for both type I and heterotic theories [18]. However, in spacetimes with a single Killing circle—effectively reducing the theory to 9 dimensions—[30] argues that the anomaly cancellation term vanishes in the heterotic theory, along with the coupling ${\rm Tr}(F^{4})$. As a result, in such backgrounds, the coupling ${\rm Tr}(F^{4})$ in type I theory can no longer be considered exact and, consequently, does not appear in the 9-dimensional heterotic theory.
@@ 3 Six-dimensional Heterotic/Type IIA duality
The sphere-level NS-NS couplings of type II have been identified in [26] by imposing T-duality on the minimal basis of NS-NS fields, which includes 872 couplings. T-duality uniquely determines all coupling constants up to one overall factor [26]. These couplings can be expressed in various schemes. In particular, the scheme in which the dilaton appears only through the overall factor $e^{-2\Phi}$ [49] includes the following terms, which contain the Riemann-squared term:
@@
where the dots refer to all other terms that do not include the Riemann-squared term $R_{\mu\nu\rho\tau}R^{\mu\nu\rho\tau}$. Up to a factor of $-2/3$, the couplings in the scheme [26] also include the aforementioned terms. Since $K3$ surfaces are Ricci-flat, the Ricci tensor vanishes. However, the Riemann curvature tensor does not. The integral of the Riemann-squared term on the $K3$ manifold is a constant, which is given by [46]
@@ 4 Discussion
In this paper, we have demonstrated that the consistency of heterotic supergravity (4) with T-duality requires an infinite number of new covariant couplings at all orders of $\alpha^{\prime}$. The couplings at orders $\alpha^{\prime}$ and $\alpha^{\prime 2}$ have already been identified in [24, 25]. This T-dual multiplet may correspond, through suitable non-covariant field redefinitions, to the locally Lorentz-invariant supersymmetric multiplet proposed in [31, 32]. Such field redefinitions at order $\alpha^{\prime}$ for NS-NS fields have been explicitly derived in [52]. Because the T-dual multiplet incorporates the leading-order two-derivative couplings, its genus dependence is exact and remains free from quantum corrections in spacetimes with a single Killing self-dual circle. However, at the classical level, there exist infinitely many additional T-dual multiplets at orders $\alpha^{\prime 3}$ and higher that do not contain the leading-order couplings [38]. Consequently, in such spacetimes, all of these additional multiplets may be subject to quantum corrections.
To clarify the final point, we note that classical couplings are background-independent. As a result, the radius of the circle in the T-duality constraint can be arbitrary, and classical T-dual multiplets in heterotic theory remain valid for any radius—even in globally flat Minkowski spacetime. In contrast, quantum corrections are background-dependent. The constraints imposed by T-duality at the quantum level generally depend on the radius of the circle. However, at the self-dual radius, the quantum T-duality constraints align with their classical counterparts [30]. Therefore, classical couplings can be extended to include quantum corrections at the self-dual radius by promoting the overall factor of the classical T-dual multiplet to one that incorporates quantum effects. This factor can be computed using the S-matrix method in backgrounds with a single self-dual circle. Importantly, the quantum T-dual multiplet that includes the leading two-derivative action receives no quantum corrections, as dictated by S-matrix calculations, which show that the leading-order effective action remains uncorrected [8, 35]. On the other hand, classical T-dual multiplets with coefficients such as $\zeta(3)$ do not contain two-derivative terms and thus do receive quantum corrections. Consequently, only the classical T-dual multiplet that includes the two-derivative couplings is exact and unaffected by quantum corrections at the self-dual radius. All other multiplets, especially those with coefficients involving $\zeta(3)$ and higher $\zeta$-functions, are subject to quantum corrections and therefore do not constitute exact T-dual multiplets. As a result, they cannot be reliably analyzed within the framework of S-duality.
We then examine the exact T-dual multiplet under the S-duality of 10-dimensional heterotic/type I theory to explore its manifestation in type I theory. These multiplets appear at various genera in type I theory. Since the T-dual multiplet involves higher-derivative couplings, their transformations under S-duality require understanding whether S-duality receives higher-derivative corrections. We observe that S-duality does not receive higher-derivative corrections. Using this insight, we identify the order $\alpha^{\prime}$ couplings in Type I theory (17), which contain no dilaton derivatives. Additionally, we utilize the S-duality of six-dimensional heterotic/type IIA theory to verify the 10-dimensional NS-NS couplings of type IIA theory at order $\alpha^{\prime 3}$, which were recently identified through T-duality.
@@
If one could calculate all genus and all $\alpha^{\prime}$ couplings in the effective action (2), then this action could generate any S-matrix element in quantum gravity that is UV finite. It would produce the result for quantum gravity that ordinary heterotic string theory produces. The higher $\alpha^{\prime}$ expansion dictates the stringy nature of quantum gravity, and the genus expansion dictates the quantum corrections required for higher-derivative couplings to be consistent with quantum theory. However, the exact T-dual multiplet has no genus expansion in spacetimes with a single Killing self-dual circle. It is
[…]
where ${\cal L}_{\rm exact}(\alpha^{\prime})$ represents the supergravity (4) and all higher-derivative exact-corrections that can, in principle, be calculated via T-duality. The higher-derivative terms highlight the stringy nature of gravity; however, each higher-derivative term lacks quantum correction. The above action represents heterotic supergravity, extended to maintain consistency with T-duality. While it is known that heterotic supergravity is not UV finite, the inclusion of an infinite number of higher-derivative couplings, akin to ordinary string theory, suggests it may achieve UV finiteness. If so, this T-dual heterotic supergravity could be proposed as a definition of quantum gravity. Given its consistency with T-duality, one might identify a particular chiral string theory, such as those proposed in [53, 54], whose low-energy effective action is (27). While this effective action has only classical couplings in one setting, its type I setting includes both classical and quantum couplings. Investigating this proposal for quantum gravity would be intriguing.

### v5

@@
This study examines the consistency of heterotic supergravity under T-duality when the $B$-field gauge transformation is rendered anomalous by the Green-Schwarz mechanism. We demonstrate that T-duality invariance mandates an infinite tower of higher-derivative couplings, all scaling as $e^{-2\Phi}$. Within this tower, the couplings at orders $\alpha^{\prime}$ and $\alpha^{\prime 2}$ are protected from quantum corrections, making them exact and therefore amenable to analysis under the S-duality between heterotic and type I string theory. Our results confirm that the standard S-duality map itself does not receive higher-derivative modifications. Leveraging this exact correspondence, we derive the explicit form of the type I effective action at order $\alpha^{\prime}$ in a scheme that omits dilaton derivatives.
@@ 1 Introduction
However, while classical couplings are background-independent, their quantum corrections are not [30]. This distinction is essential: the background independence of the classical action allows T-duality to fix its couplings in ten dimensions, whereas couplings at the quantum level remain unconstrained by this duality. Specifically, in the quantum theory, T-duality constrains couplings only in the base space after compactifying one dimension on a circle, provided all fields are independent of that direction. The background dependence at the quantum level does not allow us to use these constraints to determine the couplings in the uncompactified theory. Despite this general limitation, the classical NS-NS couplings in heterotic theory at orders $\alpha^{\prime}$ and $\alpha^{\prime 2}$ are protected from quantum corrections [8, 35]. As these couplings are quantum-exact, their characteristic genus dependence, $e^{-2\Phi}$, is also exact. This protection makes them particularly well-suited for analysis under S-duality.
@@
$\textbf{ S}_{\rm per.}$ $\sim$ $\sum_{g=0}^{\infty}\int d^{10}x\,e^{2(g-1)\Phi^{\prime}}\sqrt{-G^{\prime}}\,{\cal L}_{g}(\alpha^{\prime})\,$, (2)
@@
and no scale transformation applied to the frame $e_{\mu}{}^{\mu_{1}}$. Consequently, the spin connection $\omega_{\mu\mu_{1}}{}^{\nu_{1}}$, like the YM connection $A_{\mu i}{}^{j}$, remains invariant. As a result, the Lorentz Chern-Simons three-form $\Omega_{\mu\nu\rho}$ also remains invariant under S-duality, i.e.,
@@
This replacement is mandated by the Green-Schwarz mechanism [15], ensuring that T-duality consistently determines the coefficients of an infinite set of higher-derivative couplings at the classical level. These couplings extend the two-derivative terms in (4) to include T-dual-invariant terms at orders $\alpha^{\prime 0}$, $\alpha^{\prime}$, $\alpha^{\prime 2}$, and beyond, all governed by the same overall factor as the leading-order action [^3] A similar set of couplings arises in bosonic string theory. Specifically, bosonic theory includes an infinite number of higher-derivative couplings at orders $\alpha^{\prime}$, $\alpha^{\prime 2}$, and beyond, governed by a single overall factor [37].. The T-dual couplings at orders $\alpha^{\prime 0}$, $\alpha^{\prime}$, and $\alpha^{\prime 2}$ do not receive higher-genus corrections due to kinematic reasons [8, 35], whereas all other couplings may receive such corrections. Consequently, the couplings at these specific orders are exact and can be meaningfully studied under S-duality.
@@
Furthermore, the classical theory includes another distinct set of T-dual couplings at orders $\alpha^{\prime 3}$, $\alpha^{\prime 4}$, and higher, characterized by an overall coefficient of $\zeta(3)$. Unlike the first set, this group is not directly associated with the leading-order action (4). Similarly, there exist additional sets with coefficients $\zeta(5)$ and higher [38], all of which may receive higher-genus corrections.
@@
A similar replacement to (10) exists in Type I theory. However, the corresponding $\alpha^{\prime}H\Omega$ and $\alpha^{\prime 2}\Omega\Omega$ terms do not appear in the usual untwisted sector. Instead, they emerge in the twisted sector of the type I effective action [39], which transforms under T-duality to the twisted sector of the type I’ theory [40]. Unlike in heterotic theory, T-duality alone is insufficient to fully determine all couplings at orders $\alpha^{\prime}$, $\alpha^{\prime 2}$, and higher in both type I and type I’ theories due to inherent dimensional differences: the twisted sector in type I’ theory is inherently 9-dimensional, whereas that in type I theory is 10-dimensional [39]. Nevertheless, S-duality applied to the exact couplings of heterotic theory offers a framework to identify such couplings in type I theory.
@@
The corresponding couplings in type I theory manifest at different genera. Using the S-duality transformation (3), it can be shown that the exact couplings at order $\alpha^{\prime n}$ in heterotic theory transform into couplings at genus $e^{(n+k-2)\Phi}$ in type I theory, where $k$ represents half the number of YM field strengths in the couplings. In this paper, we demonstrate how these couplings can be systematically determined by imposing S-duality on the exact couplings in heterotic theory. Specifically, we explicitly derive these couplings in type I theory at the $\alpha^{\prime}$ order, adhering to a scheme that excludes dilaton derivatives.
@@
The exact couplings in heterotic theory has significant implications for the type IIA effective action. These implications arise from the application of S-duality between heterotic theory on the manifold $T^{4}$, which includes the circle, and type IIA theory on the manifold $K3$ [42, 16]. The S-duality correspondence between the 6-dimensional massless NS-NS fields in type IIA and their equivalents in heterotic theory is outlined in [43]:
@@
where $\Phi,G_{\mu\nu},H_{\mu\nu\rho}$ are the six-dimensional NS-NS fields of type IIA theory. In this context, one can use the dimensional reduction of the exact couplings on the torus $T^{4}$ to determine the corresponding six-dimensional couplings with the exact genus dependence $e^{-2\Phi^{\prime}}$. Using the aforementioned S-duality transformations, one can identify the corresponding higher-derivative six-dimensional couplings in type IIA theory, which appear at different genera. Applying the S-duality transformation to the exact couplings, one finds that the couplings at order $\alpha^{\prime n}$ on $T^{4}$ appear at genus $e^{2(n-1)\Phi}$ in type IIA theory on $K3$. For example, the one-loop couplings in type IIA string theory on $K3$ arise at order $\alpha^{\prime}$. These couplings were first identified in [43] through S-duality transformations and subsequently confirmed by S-matrix calculations. Furthermore, they have been utilized in [44, 45] to investigate $\alpha^{\prime}$ corrections to the dyonic string solution in 6-dimensional supergravity. The two-loop couplings in type IIA on $K3$ are at order $\alpha^{\prime 2}$, which may be found by imposing S-duality on the $\alpha^{\prime 2}$ couplings identified in [25] through T-duality.
@@
On the other hand, the aforementioned S-duality can be utilized to establish a no-go theorem. The NS-NS couplings in the heterotic theory at orders $\alpha^{\prime 0}$, $\alpha^{\prime}$, and $\alpha^{\prime 2}$ remain free from quantum corrections [8, 35]. Since the $T^{4}$ reduction of these classical couplings preserves the number of derivatives in the couplings, it follows that no couplings arise at orders $\alpha^{\prime}$ and $\alpha^{\prime 2}$ at the quantum level, i.e., at genus $e^{2(g-1)\Phi^{\prime}}$ for $g>0$, in the six-dimensional heterotic theory. Using the S-duality transformation, one finds no couplings at order $\alpha^{\prime}$ at genus $e^{-2g\Phi}$ and no couplings at order $\alpha^{\prime 2}$ at genus $e^{-2(g-1)\Phi}$ in the six-dimensional type IIA theory. Based on the permitted genus expansion for the effective action in type IIA theory, it is determined that at order $\alpha^{\prime}$, $g=1$, and at order $\alpha^{\prime 2}$, $g=1,2$. Consequently, there must be no classical couplings at order $\alpha^{\prime}$, and no classical or torus-level couplings at order $\alpha^{\prime 2}$ in the six-dimensional type IIA theory. This conclusion is reinforced by arguments derived from supersymmetry (see, for example, [46]). We will demonstrate that the $K3$ reduction of the classical couplings in type IIA theory at order $\alpha^{\prime 3}$, identified via T-duality [26], produces certain $\alpha^{\prime}$ couplings that vanish upon applying field redefinition.
@@ 2 Heterotic/Type I duality
$\Delta B_{\alpha\beta}(\Phi,G,B)+\Delta B_{\alpha\beta}(-\Phi,e^{-\Phi}G,B)$ $=$ $0.$ (13)
@@
$+\frac{1}{48}F_{\alpha\beta ij}F^{\alpha\beta ij}{\tilde{H}}_{\gamma\delta\epsilon}{\tilde{H}}^{\gamma\delta\epsilon}-\frac{1}{4}F^{\alpha\beta ij}{\tilde{H}}_{\beta\gamma\delta}\nabla_{\alpha}F^{\gamma\delta}{}_{ij}+\frac{3}{64}F_{\alpha\beta ij}F^{\alpha\beta ij}R-\frac{1}{2}F_{\alpha}{}^{\gamma ij}F_{\beta\gamma ij}R^{\alpha\beta}\Big]$
$+e^{-\Phi}\Big[\frac{1}{16}R^{2}-\frac{1}{16}{\tilde{H}}_{\alpha\beta}{}^{\delta}{\tilde{H}}^{\alpha\beta\gamma}{\tilde{H}}_{\gamma}{}^{\epsilon\varepsilon}{\tilde{H}}_{\delta\epsilon\varepsilon}+\frac{1}{96}{\tilde{H}}_{\alpha\beta\gamma}{\tilde{H}}^{\alpha\beta\gamma}{\tilde{H}}_{\delta\epsilon\varepsilon}{\tilde{H}}^{\delta\epsilon\varepsilon}-R_{\alpha\beta}R^{\alpha\beta}$
@@
where ${\tilde{H}}=e^{\Phi}H$. Note that the YM couplings in the first bracket are double traces. The couplings in the last bracket occur at the disk level, while the others appear at higher genus levels. At the four-derivative order, the disk-level effective action of type I theory incorporates the single-trace coupling ${\rm Tr}(F^{4})$ as well. In heterotic theory, this coupling emerges as a one-loop contribution and is related, via a supersymmetry transformation, to the anomaly cancellation mechanism in 10-dimensional spacetime. Consequently, it represents an exact coupling in 10 dimensions for both type I and heterotic theories [18].
@@ 3 Six-dimensional Heterotic/Type IIA duality
The sphere-level NS-NS couplings of type II at order $\alpha^{\prime 3}$ have been identified in [26] by imposing T-duality on the minimal basis of NS-NS fields, which includes 872 couplings. T-duality uniquely determines all coupling constants up to one overall factor [26]. These couplings can be expressed in various schemes. In particular, the scheme in which the dilaton appears only through the overall factor $e^{-2\Phi}$ [49] includes the following terms, which contain the Riemann-squared term:
@@
where the dots refer to all other terms that do not include the Riemann-squared term $R_{\mu\nu\rho\tau}R^{\mu\nu\rho\tau}$. Up to a factor of 2/3-2/3, the couplings in the scheme [26] also include the aforementioned terms. Since $K3$ surfaces are Ricci-flat, the Ricci tensor vanishes. However, the Riemann curvature tensor does not. The integral of the Riemann-squared term on the $K3$ manifold is a constant, which is given by [46]
@@ 4 Discussion
In this paper, we have demonstrated that the consistency of heterotic supergravity (4) under T-duality requires an infinite tower of covariant higher-derivative couplings at all orders in $\alpha^{\prime}$. The explicit forms of the couplings at orders $\alpha^{\prime}$ and $\alpha^{\prime 2}$ have been identified in previous work [24, 25]. This T-dual-invariant multiplet can be related, via appropriate non-covariant field redefinitions, to the locally Lorentz-invariant supersymmetric multiplet proposed in [31, 32]. Such field redefinitions for the NS-NS sector at order $\alpha^{\prime}$ have been explicitly constructed in [52]. At the classical level, there also exist infinitely many distinct T-dual multiplets at orders $\alpha^{\prime 3}$ and above, which do not include the leading two-derivative terms [38]. Crucially, the $\alpha^{\prime}$ and $\alpha^{\prime 2}$ couplings within the T-dual multiplet are protected from quantum corrections, as established by S-matrix calculations [8, 35]. In contrast, all higher-order couplings are subject to quantum corrections and thus cannot be studied reliably under S-duality.
We then examine the exact couplings in the context of S-duality between 10-dimensional heterotic and type I theories to determine their manifestation in type I. In type I, these couplings appear across various genera. As they are higher-derivative terms, their behavior under S-duality hinges on whether the S-duality map itself receives higher-derivative corrections. We observe that S-duality does not receive such corrections. Utilizing this insight, we identify the corresponding order $\alpha^{\prime}$ couplings in type I theory (17), which contain no dilaton derivatives. Furthermore, we employ the S-duality between six-dimensional heterotic and type IIA theories to verify the 10-dimensional NS-NS couplings of type IIA theory at order $\alpha^{\prime 3}$, which were recently identified via T-duality.